# Complementary Time–Space Tradeoff for Self-Stabilizing Leader Election: Polynomial States Meet Sublinear Time

### v1

@@ Abstract
We study the self-stabilizing leader election (SS-LE) problem in the population protocol model, assuming exact knowledge of the population size $n$. Burman, Chen, Chen, Doty, Nowak, Severson, and Xu (PODC 2021) showed that this problem can be solved in $O(n)$ expected time with $O(n)$ states. Recently, Gąsieniec, Grodzicki, and Stachowiak (PODC 2025) proved that $n+O(\log n)$ states suffice to achieve $O(n\log n)$ time both in expectation and with high probability (w.h.p.). If substantially more states are available, sublinear time can be achieved. Burman et al. (PODC 2021) presented a $2^{O(n^{\rho}\log n)}$-state SS-LE protocol with a parameter $\rho$: setting $\rho=\Theta(\log n)$ yields an optimal $O(\log n)$ time both in expectation and w.h.p., while $\rho=\Theta(1)$ results in $O(\rho\,n^{1/(\rho+1)})$ expected time. Very recently, Austin, Berenbrink, Friedetzky, Götte, and Hintze (PODC 2025) presented a novel SS-LE protocol parameterized by a positive integer $\rho$ with $21\leq\rho<n/$ 2 that solves SS-LE in $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ time w.h.p. using $2^{O(\rho^{2}\log n)}$ states. This paper independently presents yet another time–space tradeoff of SS-LE: for any positive integer $\rho$ with $1\leq\rho\leq\sqrt{n}$, SS-LE can be achieved within $O\left(\nicefrac{{n}}{{\rho}}\cdot\log\rho\right)$ expected time using $2^{2\rho\lg\rho+O(\log n)}$ states. The proposed protocol uses significantly fewer states than the protocol of Austin et al. requires to achieve any expected stabilization time above $\Theta(\sqrt{n}\log n)$. When $\rho=\Theta\left(\nicefrac{{\log n}}{{\log\log n}}\right)$, the proposed protocol is the first to achieve sublinear time while using only polynomially many states. A limitation of our protocol is that the constraint $\rho\leq\sqrt{n}$ prevents achieving $o(\sqrt{n}\log n)$ time, whereas the protocol of Austin et al. can surpass this bound.
@@ 1 Introduction
This paper investigates the self-stabilizing leader election problem (SS-LE) in the population protocol model introduced by Angluin, Aspnes, Diamadi, Fischer, and Peralta [6], which has been extensively studied in distributed computing for over two decades. Recently, the time–space tradeoff for SS-LE under exact knowledge of the population size $n$ has attracted intensive study, particularly in 2025. We present a novel time–space tradeoff that dramatically improves the upper bound on space complexity required to achieve any target time complexity in the range between $\Theta(\sqrt{n}\log n)$ and $o(n)$.
@@ 1.1 Population Protocols
Formally, a protocol is a five-tuple $\mathcal{P}=(Q,s_{0},Y,\delta,\pi)$, where $Q$ is the set of agent states, $s_{0}$ is the initial state (unused in self-stabilizing protocols), $Y$ is the set of output symbols, $\delta\colon Q\times Q\to Q\times Q$ is the transition function, and $\pi\colon Q\to Y$ is the output function.111 We omit the input symbol set and the input function here, as the leader election problem involves no external inputs. A global state of the population (or configuration) is a function $C\colon A\to Q$ that represents the current state of each agent. A configuration $C$ is output-stable (or simply stable) if no agent ever changes its output in any subsequent execution from $C$.
The complexity of a protocol is measured by its stabilization time and number of states. The stabilization time of a protocol $\mathcal{P}$ is the number of interactions divided by $n$ required for the population under $\mathcal{P}$ to reach a stable configuration. [^2] This definition applies to static problems such as leader election and majority; a more intricate definition is required for non-static problems. See [5]. Generally, the number of interactions divided by $n$ is called parallel time, reflecting the practical behavior where agents often interact concurrently. Throughout this paper, “time” refers to parallel time. Since interacting agents are chosen uniformly at random, we evaluate stabilization time in expectation or with high probability. We define “with high probability”(w.h.p.) as an event occurring with probability $1-O(n^{-\eta})$ for any fixed positive constant $\eta=\Theta(1)$ (see Section 2.1 for the formal definition). The number of states of a protocol is defined as the cardinality of its state set $Q$, i.e., $|Q|$. Because super-exponentially many states are sometimes used, we also express space complexity in bits: $x$ bits correspond to $2^{x}$ states.
@@ 1.2 Related Work
Time–space tradeoffs arise for many problems: the more states available, the shorter the stabilization time. For example, the time–space tradeoff of the (non-self-stabilizing) leader election problem was intensively studied from 2015 to 2020. As the name implies, the leader election problem requires exactly one agent to output ‘L’ (leader), while all others output ‘F’ (follower). Doty and Soloveichik [20] proved that any $O(1)$-state protocol requires $\Omega(n)$ expected time to solve leader election (DISC 2015), establishing the time-optimality of the two-state protocol by Angluin et al. [6], which stabilizes in $O(n)$ expected time. Simultaneously, Alistarh and Gelashvili [4] presented a protocol using $O(\log^{3}n)$ states and achieving $O(\log^{3}n)$ expected time (ICALP 2015). Subsequent work [1, 13, 2, 23, 24, 31] successively improved these upper bounds from 2017 to 2019. Finally, Berenbrink, Giakkoupis, and Kling [12] presented an $O(\log\log n)$-state and $O(\log n)$-expected-time protocol (STOC 2020). This protocol is both time- and space-optimal, owing to two lower bounds: Alistarh, Aspnes, Eisenstat, Gelashvili, and Rivest [1] showed that any protocol with fewer than $\tfrac{1}{2}\log\log n$ states requires $\Omega(n^{2}/\operatorname{poly}(\log n))$ expected time; and Sudo and Masuzawa [29] proved that any protocol requires $\Omega(\log n)$ states. [^3] The latter lower bound goes beyond the simple coupon-collector argument, as it must handle initial configurations with all agents in the follower state.
@@
As mentioned above, this paper addresses the self-stabilizing leader election problem (SS-LE). Self-stabilization, introduced by Dijkstra [19], is a fault-tolerance concept that requires a protocol to converge to a correct configuration even when execution begins from an arbitrary configuration. A population protocol $\mathcal{P}$ solves SS-LE if, from any initial configuration, it eventually stabilizes with exactly one leader. Solving SS-LE is more challenging than (non-self-stabilizing) leader election. An SS-LE protocol must create a leader when none exists and reduce multiple leaders to one; these opposing goals often conflict. Indeed, using the partition technique of [5, 16], one can easily show that SS-LE is impossible unless each agent knows the exact population size $n$ (see [34], p. 618, footnote). Despite this impossibility result, SS-LE has been continuously studied under additional assumptions or relaxed self-stabilization requirements in the population protocol model.
@@
One approach to solving SS-LE is to assume that each agent knows the exact population size $n$, i.e., $n$ is hard-coded into the protocol. Tables 2 and 2 summarize the upper and lower bounds in this setting, including our contributions, for silent and non-silent SS-LE protocols, respectively. Silent protocols require the population to eventually reach a configuration from which no agent changes its state in any subsequent execution, a stronger requirement than stabilization. In some protocols listed in these tables, the expected and w.h.p. stabilization times coincide; in others, they differ. Nonetheless, the discrepancy is at most an $O(\log n)$ multiplicative factor. By the definition of self-stabilizing protocols, the following relationship holds:
@@
For any function $f:\mathbb{N}^{+}\to\mathbb{N}^{+}$, if a protocol solves SS-LE in $O(f(n))$ expected time, then it solves SS-LE in $O(f(n)\log n)$ time w.h.p. Conversely, if a protocol solves SS-LE in $O(f(n))$ time w.h.p., then it solves SS-LE in $O(f(n))$ expected time.
It is also worth mentioning here that all SS-LE protocols that assume exact knowledge of the population size $n$, listed in Tables 2 and 2, also solve the more general problem of self-stabilizing ranking (SS-RK). This problem requires $n$ agents, starting from an arbitrary configuration, to stabilize to have distinct labels (or ranks) in $\{1,2,\dots,n\}$. Once they obtain unique ranks, the agent with rank $1$ can serve as the unique leader; hence, SS-LE reduces to SS-RK, while the converse reduction remains open. Nevertheless, the fact that all existing SS-LE protocols (assuming only knowledge of $n$) solve SS-RK suggests the following conjecture.
@@
Every SS-LE protocol solves SS-RK, at least implicitly: for any SS-LE protocol $\mathcal{P}$ with state set $Q$, there exists a function $\xi\colon Q\to\{1,2,\dots,n\}$ such that for every stable configuration $C$ of $\mathcal{P}$, |$|\{\xi(C(a)):a\in A\}|=n$.
Assuming knowledge of $n$, Cai, Izumi, and Wada [16] showed that any SS-LE protocol requires at least $n$ states and presented a silent, state-optimal ($n$-state), $O(n^{2})$-expected-time SS-LE protocol. Later, Burman, Chen, Chen, Doty, Nowak, Severson, and Xu [14] showed that increasing the number of states by a constant factor reduces the expected stabilization time to $O(n)$. They also proved this bound is optimal: any silent SS-LE protocol requires $\Omega(n)$ expected stabilization time (and $\Omega(n\log n)$ time w.h.p.). Dropping the silent requirement enables sublinear time (i.e., $o(n)$ time): they presented an SS-LE protocol that stabilizes in $O(\log n)$ time w.h.p., using super-exponentially many states (i.e., $O(n^{\log n}\cdot\log n)$ bits). No faster protocol exists, since any SS-LE protocol requires $\Omega(\log n)$ time—a trivial lower bound following from the coupon collector argument, unlike the non-self-stabilizing case [29]. They also proposed a protocol parameterized by a constant $\rho=\Theta(1)$ that stabilizes in $O(\rho\,n^{1/(\rho+1)})$ expected time using a reduced yet still super-exponential number of states, namely $O(n^{\rho}\log n)$ bits.
Table 1: Silent SS-LE protocols with exact knowledge of the population size $n$ (space shown in number of states); all protocols also solve SS-RK.
@@
[16] $O(n^{2})$ $O(n^{2})$ $n$
[14] $O(n)$ $O(n\log n)$ $O(n)$
[11] $O(n\log n)$ $O(n\log n)$ $n+O(\log^{2}n)$
[25] $O(n\log n)$ $O(n\log n)$ $n+O(\log n)$
[25] $O(n^{7/4}\log^{2}n)$ $O(n^{7/4}\log^{2}n)$ $n+1$
[16] any any $\geq n$
[14] $\Omega(n)$ $\Omega(n\log n)$ any
Table 2: Non-silent SS-LE protocols with exact knowledge of the population size $n$ (space shown in bits; $x$ bits correspond to $2^{x}$ states); all protocols also solve SS-RK.
@@
[14] $O(\log n)$ $O(\log n)$ $O(n^{\log n}\log n)$
[14] $O(\rho\cdot n^{1/(\rho+1)})$ $O(\rho\cdot n^{1/(\rho+1)}\log n)$ $O(n^{\rho}\log n)$ $\rho\geq 1,\rho=\Theta(1)$
[9] $O(\frac{n}{\rho}\log n)$ $O(\frac{n}{\rho}\log n)$ $O(\rho^{2}\log n)$ $21\leq\rho<n/$ 2
Theorem 1 $O(\frac{n}{\rho}\log\rho)$ $O(\frac{n}{\rho}\log\rho\cdot\log n)$ $2\rho\lg\rho+O(\log n)$ $1\leq\rho\leq\sqrt{n}$
($\rho=\sqrt{n}$) $O(\sqrt{n}\log n)$ $O(\sqrt{n}\log^{2}n)$ $\frac{\sqrt{n}\lg n}{2}+O(\log n)$
($\rho=\frac{\log n}{\log\log n}$) $O(\frac{n\log^{2}\log n}{\log n})$ $O(n\log^{2}\log n)$ $O(\log n)$
@@
Building on Burman et al. [14], SS-LE protocols with exact knowledge of $n$ have attracted intensive study recently. Four new works—Berenbrink, Elsässer, Götte, Hintze, and Kaaser [11]; Gąsieniec, Grodzicki, and Stachowiak [25]; Araya and Sudo [8]; and Austin, Berenbrink, Friedetzky, Götte, and Hintze [9]— were made public on arXiv in the first half of 2025 and will appear in the proceedings of ICDCS 2025, PODC 2025, SIROCCO 2025, and PODC 2025, respectively.
The first two studies, Berenbrink et al. [11] and Gąsieniec et al. [25], provide new insights on silent SS-LE. Recall that any SS-LE protocol requires at least $n$ states. A natural question is how many additional states suffice to substantially reduce the stabilization time of the space-optimal protocol of Cai et al. [16], which stabilizes in $O(n^{2})$ time (both in expectation and w.h.p.). Berenbrink et al. [11] showed that adding $O(\log^{2}n)$ states yields a silent SS-LE protocol within optimal $O(n\log n)$ stabilization time w.h.p. Gąsieniec et al. [25] further reduced the additional states to $O(\log n)$ while preserving optimal w.h.p. time. Moreover, they proved that a single additional state suffices to achieve $O(n^{7/4}\log^{2}n)$ time w.h.p., still maintaining the silent property.
For non-silent SS-LE research, Burman et al. [14] identified promising directions. They proved that dropping the silent requirement enables sublinear stabilization time. However, all their sublinear-time protocols require a super-exponential number of states (i.e., $\omega(n)$ bits), even when minimizing space by setting $\rho=1$, which yields an $O(\sqrt{n})$-expected-time protocol. Moreover, if Conjecture 1 holds, SS-LE has the same time–space tradeoff as SS-RK, which inherently requires detecting rank collisions (i.e., two or more agents with the same rank). In their words, “the core difficulty” of fast SS-RK lies in collision detection. Indeed, their non-silent protocols rely on a super-exponential number of states for this task. Based on this discussion, Burman et al. [14] posed the following open problems:
@@
    Does there exist an SS-LE protocol that achieves sublinear stabilization time using only a subexponential number of states (i.e., $o(n)$ bits, which results in $o(c^{n})$ states for any constant $c>1$)?
@@
    Does there exist a protocol that detects collisions in sublinear time using only a subexponential number of states even in the non-self-stabilizing setting? Here, “non-self-stabilizing” means that agents begin with adversarially assigned ranks but otherwise share the designated initial state.
@@
Araya and Sudo [8] answered the second question affirmatively: a near-linear number of states suffices to detect collisions in polynomial time. Specifically, they proposed a protocol that, in the non-self-stabilizing setting, detects collisions in $O(\sqrt{n\log n})$ expected time using only $O(n\cdot\mathrm{poly}(\log n))$ states. However, this protocol does not immediately resolve the “core difficulty” of fast SS-RK. It comprises two sub-protocols—a fast primary protocol and a slow backup protocol—executed in parallel by each agent. The primary protocol stabilizes in $O(\sqrt{n}\,\log^{3/2}n)$ time w.h.p. but eventually terminates, ceasing collision-detection attempts; this termination is necessary to avoid false positives. The backup protocol never terminates but requires $\Theta(n)$ time. Together, they detect any collision with probability $1$ in $O(\sqrt{n\log n})$-expected time. (See [8] for the detailed analysis.) To apply this protocol to SS-RK, we must repeatedly invoke the primary protocol and reset its internal state across all agents. Without this reset, execution may start in an arbitrary configuration where the primary protocol has already terminated; in that case, only the backup protocol can detect collisions, incurring an inevitable $\Theta(n)$ delay. However, repeating this collision-detection cycle also introduces a critical issue: since a pair of agents is chosen uniformly at random (we write u.a.r. in the rest of this paper) at each time step, there remains a nonzero (albeit arbitrarily small) probability that a global reset fails—some agents may retain outdated information even when execution begins from a stable configuration. This residual information causes false-positive detections and forces rank reassignment, thereby violating stability.
@@
Very recently, Austin et al. [9] presented a dramatically improved time–space tradeoff for SS-LE, affirmatively answering the first open question. They proposed a protocol parameterized by a positive integer $\rho$ ($21\leq\rho<n/$ 2) that solves SS-RK (and thus SS-LE) in $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ time w.h.p. using $O(\rho^{2}\log n)$ bits. Choosing $\rho=\omega(\log n)\cap o(\sqrt{n/\log n})$ yields a sublinear-time, sub-exponential-state protocol: for example, $\rho=\Theta(\sqrt{n}/\log n)$ gives $O(\sqrt{n}\log^{2}n)$ time w.h.p. with $2^{O(n/\log n)}=2^{o(n)}$ states, while $\rho=\Theta(\log^{2}n)$ gives $O(n/\log n)$ time w.h.p. with $2^{O(\log^{5}n)}$ states. Allowing super-exponentially many states—but still far fewer than in [14]—yields a time-optimal solution: setting $\rho=\Theta(n)$ achieves $\Theta(\log n)$ time w.h.p. using $2^{O(n^{2}\log n)}$ states. Their collision-detection module also independently resolves the second question posed to Araya and Sudo [8]. Although it uses significantly more states ($2^{O(\rho^{2}\log n)}$ vs. $O(n\cdot\mathrm{poly}(\log n))$), it never terminates and thus does not require repeated invocation, enabling the first sublinear-time, sub-exponential-state SS-RK protocol.
@@
While the majority work on population protocols studies the setting where every pair of agents can interact, quite a few studies deal with a more general setting: interactions may occur between restricted pairs of agents, introducing a interaction graph $G=(V,E)$, where $V$ represents the set of agents and $E$ the set of interactable pairs. SS-LE and its weaken variant have been also studied in rings [21, 17, 36, 37], regular graphs [18], and general graphs [21, 33, 32, 35, 26].

### v2

# arXiv:2505.23649v2[cs.DC] 27 Feb 2026

@@ Abstract
We study the self-stabilizing leader election (SS-LE) problem in the population protocol model, assuming exact knowledge of the population size $n$. Burman, Chen, Chen, Doty, Nowak, Severson, and Xu [BCC+21] (PODC) showed that this problem can be solved in $O(n)$ expected time with $O(n)$ states. Recently, Gąsieniec, Grodzicki, and Stachowiak [GGS25] (PODC) proved that $n+O(\log n)$ states suffice to achieve $O(n\log n)$ time both in expectation and with high probability (w.h.p.). If substantially more states are available, sublinear time can be achieved. The authors of [BCC+21] presented a $2^{O(n^{\rho}\log n)}$-state SS-LE protocol with a parameter $\rho$: setting $\rho=\Theta(\log n)$ yields an optimal $O(\log n)$ time both in expectation and w.h.p., while $\rho=\Theta(1)$ results in $O(\rho\,n^{1/(\rho+1)})$ expected time. Recently, Austin, Berenbrink, Friedetzky, Götte, and Hintze [ABF+25] (PODC) presented a novel SS-LE protocol parameterized by a positive integer $\rho$ with $1\leq\rho<n/2$ that solves SS-LE in $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ time w.h.p. using $2^{O(\rho^{2}\log n)}$ states. This paper independently presents yet another time–space tradeoff of SS-LE: for any positive integer $\rho$ with $2\leq\rho\leq\sqrt{n}$, SS-LE can be achieved within $O\left(\nicefrac{{n}}{{\rho}}\cdot\log\rho\right)$ expected time using $2^{2\rho\lg^{2}\rho+O(\log n)}$ states. The proposed protocol uses significantly fewer states than [ABF+25] for any expected stabilization time above $\Theta(\sqrt{n}\log n)$. When $\rho=\Theta\left(\nicefrac{{\log n}}{{\log^{2}\log n}}\right)$, the proposed protocol is the first to achieve sublinear time while using only polynomially many states. A limitation of our protocol is that the constraint $\rho\leq\sqrt{n}$ prevents achieving $o(\sqrt{n}\log n)$ time, whereas the protocol of [ABF+25] can surpass this bound.
@@ 1 Introduction
This paper investigates the self-stabilizing leader election (SS-LE) problem in the population protocol model introduced by Angluin, Aspnes, Diamadi, Fischer, and Peralta [AAD+06], which has been extensively studied in the distributed computing community for over two decades. Recently, the time–space tradeoff for SS-LE, assuming exact knowledge of the population size $n$, has drawn significant attention. We present a novel time–space tradeoff that substantially improves the best-known upper bound on space complexity required to achieve any target time complexity within the range $\Theta(\sqrt{n}\log n)$ to $o(n)$.
@@ 1.1 Population Protocols
A protocol is a five-tuple $\mathcal{P}=(Q,s_{0},Y,\delta,\pi)$, where $Q$ is the set of agent states, $s_{0}$ is the initial state (unused in self-stabilizing protocols that we will see later), $Y$ is the set of output symbols, $\delta\colon Q\times Q\to Q\times Q$ is the transition function, and $\pi\colon Q\to Y$ is the output function. [^1] We omit the input symbol set and the input function here, as the leader election involves no external inputs. A global state of the population (or configuration) is a function $C\colon A\to Q$ that represents the state of each agent. A configuration $C$ is output-stable (or simply stable) if no agent ever changes its output in any subsequent execution from $C$.
The complexity of a protocol is measured by its stabilization time and its number of states. The stabilization time of a protocol $\mathcal{P}$ is the number of interactions, divided by the population size $n=|A|$, required for the population under $\mathcal{P}$ to reach a stable configuration. [^2] This definition applies to static problems such as leader election and majority; a more intricate definition is required for non-static problems. See [AAFJ08]. The quantity “number of interactions divided by nn” is commonly called parallel time, reflecting the fact that agents may interact concurrently in practice. Throughout this paper, “time” refers to parallel time. Since interacting agents are chosen uniformly at random, we evaluate stabilization time either in expectation or with high probability. We say that an event occurs with high probability (w.h.p.) if, for any positive constant $\eta=\Theta(1)$, it occurs with probability $1-O(n^{-\eta})$ (by appropriately adjusting the constant parameters of a protocol, if necessary). The number of states of a protocol is $|Q|$, the cardinality of its state set $Q$. Because some protocols use super-exponentially many states, we also express space complexity in bits: $x$ bits correspond to $2^{x}$ states.
@@ 1.2 Related Work
Time–space tradeoffs arise for many problems in population protocols: the more states available, the shorter the stabilization time. The tradeoff for (non-self-stabilizing) leader election was intensively studied between 2015 and 2020. In this problem, exactly one agent must output ‘L’ (leader) and all others ‘F’ (follower). Doty and Soloveichik [DS18] proved that any $O(1)$-state protocol requires $\Omega(n)$ expected time, establishing the time-optimality of the two-state protocol of [AAD+06], which stabilizes in $O(n)$ expected time. Alistarh and Gelashvili [AG18] gave a protocol with $O(\log^{3}n)$ states and $O(\log^{3}n)$ expected time, and subsequent work [AAE+17, BCER17, AAG18, BKKO18, GS20, GSU19, SOI+20] progressively improved these bounds. Finally, Berenbrink, Giakkoupis, and Kling [BGK20] obtained an $O(\log\log n)$-state and $O(\log n)$-expected-time protocol, which is optimal in both time and space. Optimality follows from two lower bounds: $o(n^{2}/\operatorname{poly}(\log n))$ expected time requires at least $\tfrac{1}{2}\log\log n$ states [AAE+17], and $\Omega(\log n)$ expected time is necessary [SM20]. [^3] The latter bound goes beyond the simple coupon-collector argument, since all agents may initially be in the follower state.
@@
This paper studies self-stabilizing leader election (SS-LE). Self-stabilization, introduced by Dijkstra [Dij74], is a fault-tolerance concept that requires a protocol to converge to a correct configuration starting from an arbitrary configuration. A population protocol $\mathcal{P}$ solves SS-LE if, starting from any configuration, it eventually stabilizes with exactly one leader. Solving SS-LE is more challenging than (non-self-stabilizing) leader election. An SS-LE protocol must create a leader when none exists and reduce multiple leaders to one; these opposing goals often conflict. Indeed, using the partition technique of [AAFJ08, CIW12], one can easily show that SS-LE is impossible unless each agent knows the exact population size $n$ (see [SOK+20], p. 618). Nevertheless, SS-LE has been extensively studied under additional assumptions or relaxed stabilization requirements.
@@
One approach to SS-LE is to assume that each agent knows the exact population size $n$, i.e., $n$ is hard-coded into the protocol. Tables 2 and 2 summarize the upper and lower bounds in this setting, including our contributions, for silent and non-silent protocols. Silent protocols require the population to reach a configuration from which no agent ever changes its state, a stronger notion than output stabilization. Note that expected and w.h.p. stabilization times differ by at most an $O(\log n)$ factor. By standard arguments for self-stabilizing protocols, the following holds:
@@
For any function $f:\mathbb{N}^{+}\to\mathbb{N}^{+}$, if a protocol solves SS-LE in $O(f(n))$ expected time, then it solves SS-LE in $O(f(n)\log n)$ time w.h.p.; conversely, $O(f(n))$ time w.h.p. implies $O(f(n))$ expected time.
All existing SS-LE protocols with exact knowledge of the population size $n$ (Tables 2 and 2) also solve the more general problem of self-stabilizing ranking (SS-RK).
Definition 1 (Self-Stabilizing Ranking (SS-RK)).
A protocol $\mathcal{P}$ solves SS-RK if, starting from any configuration, the population under $\mathcal{P}$ reaches, with probability 1, an output-stable configuration in which all agents output distinct integers in $[1,n]$.
We call such outputs ranks. Once unique ranks are assigned, the agent with rank $1$ can serve as the leader; hence SS-LE reduces to SS-RK, while the converse reduction remains open. Since all known SS-LE protocols with knowledge of $n$ also solve SS-RK, this motivates the following conjecture.
@@
SS-LE and SS-RK have the same time–space tradeoffs: for any functions $f,g:\mathbb{N}\to\mathbb{N}$, SS-LE can be solved in expected (resp. w.h.p.) stabilization time $O(f(n))$ using $O(g(n))$ states if and only if SS-RK can be solved in expected (resp. w.h.p.) stabilization time $O(f(n))$ using $O(g(n))$ states.
Assuming knowledge of $n$, Cai, Izumi, and Wada [CIW12] proved that any SS-LE protocol requires at least $n$ states and presented a silent, state-optimal ($n$-state) protocol with $O(n^{2})$ expected time. Burman, Chen, Chen, Doty, Nowak, Severson, and Xu [BCC+21a] showed that increasing the number of states by a constant factor reduces the expected stabilization time to $O(n)$, and this is time-optimal for silent SS-LE: any silent protocol requires $\Omega(n)$ expected time (and $\Omega(n\log n)$ w.h.p.). Dropping the silent requirement enables sublinear time (i.e., $o(n)$ time). They presented an $O(\log n)$-time protocol w.h.p. using super-exponentially many states (i.e., $O(n^{\log n}\log n)$ bits). No faster protocol exists, which just follows from the coupon collector argument (in contrast to the non-self-stabilizing case [SM20]). They further gave a protocol parameterized by $\rho=\Theta(1)$ achieving $O(\rho\,n^{1/(\rho+1)})$ expected time using $O(n^{\rho}\log n)$ bits.
Table 1: Silent SS-LE protocols with exact knowledge of the population size $n$ (space shown in number of states); all protocols also solve $S$-RK.
@@
[CIW12] $O(n^{2})$ $O(n^{2})$ $n$
[BCC+21a] $O(n)$ $O(n\log n)$ $O(n)$
[BEG+25] $O(n\log n)$ $O(n\log n)$ $n+O(\log^{2}n)$
[GGS25] $O(n\log n)$ $O(n\log n)$ $n+O(\log n)$
[GGS25] $O(n^{7/4}\log^{2}n)$ $O(n^{7/4}\log^{2}n)$ $n+1$
[CIW12] any any $\geq n$
[BCC+21a] $\Omega(n)$ $\Omega(n\log n)$ any
Table 2: Non-silent SS-LE protocols with exact knowledge of the population size $n$ (space shown in bits; $x$ bits correspond to $2^{x}$ states); all protocols also solve $S$-RK.
@@
[BCC+21a] $O(\log n)$ $O(\log n)$ $O(n^{\log n}\log n)$ -
[BCC+21a] $O(\rho\cdot n^{1/(\rho+1)})$ $O(\rho\cdot n^{1/(\rho+1)}\log n)$ $O(n^{\rho}\log n)$ $\rho\geq 1,\rho=\Theta(1)$
[ABF+25] $O(\frac{n}{\rho}\log n)$ $O(\frac{n}{\rho}\log n)$ $O(\rho^{2}\log n)$ $1\leq\rho<n/2$
Theorem 1 $O(\frac{n}{\rho}\log\rho)$ $O(\frac{n}{\rho}\log\rho\cdot\log n)$ $2\rho\lg^{2}\rho+O(\log n)$ $2\leq\rho\leq\sqrt{n}$
($\rho=\lfloor\sqrt{n}\rfloor$) $O(\sqrt{n}\log n)$ $O(\sqrt{n}\log^{2}n)$ $\sqrt{n}\lg^{2}n+O(\log n)$ -
($\rho=\frac{\log n}{\log^{2}\log n}$) $O(\frac{n\log^{3}\log n}{\log n})=o(n)$ $O(n\log^{3}\log n)$ $O(\log n)$ -
@@
Building on the work of [BCC+21a], SS-LE with exact knowledge of $n$ have recently received intensive attention. In particular, four new papers were published in 2025: Berenbrink, Elsässer, Götte, Hintze, and Kaaser [BEG+25]; Gąsieniec, Grodzicki, and Stachowiak [GGS25]; Araya and Sudo [AS25]; and Austin, Berenbrink, Friedetzky, Götte, and Hintze [ABF+25].
The first two studies, [BEG+25] and [GGS25], advance silent SS-LE. Since any SS-LE protocol requires at least $n$ states, a central question is how many additional states reduce the $O(n^{2})$ stabilization time of the space-optimal protocol of [CIW12]. The authors of [BEG+25] showed that $n+O(\log^{2}n)$ states achieve optimal $O(n\log n)$ time w.h.p., and [GGS25] reduced the overhead to $O(\log n)$ states while preserving optimal w.h.p. time. They further proved that even one additional state yields $O(n^{7/4}\log^{2}n)$ time w.h.p., still maintaining silence.
For non-silent SS-LE, [BCC+21a] showed that dropping silence enables sublinear stabilization time, but all their such protocols require super-exponentially many states (i.e., $\omega(n)$ bits). If Conjecture 1 holds, then fast SS-RK (and hence SS-LE) necessarily involves detecting rank collisions (i.e., two or more agents sharing the same rank). In their words, “the core difficulty” of fast SS-RK lies in collision detection, and their non-silent protocols use super-exponentially many states for this task. Motivated by this, [BCC+21a] posed the following open problems:
@@
    Does there exist an SS-LE protocol with sublinear stabilization time using only $o(n)$ bits (i.e., sub-exponentially many states)?
@@
    Does there exist a protocol that detects collisions in sublinear time using only $o(n)$ bits, even in the non-self-stabilizing setting (where agents start with adversarially assigned ranks but otherwise share the designated initial state)?
@@
The second question was answered affirmatively in [AS25], which showed that $O(n\operatorname{poly}(\log n))$ states suffice to detect collisions in polynomial time. In particular, in the non-self-stabilizing setting, collisions can be detected in $O(\sqrt{n\log n})$ expected time. However, this protocol does not directly resolve the core difficulty of fast SS-RK. It runs a fast primary protocol and a slow backup protocol in parallel: the primary stabilizes in $O(\sqrt{n}\,\log^{3/2}n)$ time w.h.p. but eventually terminates to avoid false positives, whereas the backup never terminates and requires $\Theta(n)$ expected time. Although together they detect collisions in $O(\sqrt{n\log n})$ expected time, applying this approach to SS-RK requires repeatedly restarting the primary protocol. Without this reset, execution may start in a configuration where the primary protocol has already terminated; in that case, only the backup protocol can detect collisions, incurring an inevitable $\Theta(n)$ delay. However, repeating the cycle of the primary protocol is also problematic: since interactions are chosen uniformly at random (u.a.r.) at each time step, there remains a nonzero (albeit arbitrarily small) probability that a global reset fails and some agents retain outdated information, leading to false positives and repeated rank reassignment, thereby violating stability.
@@
Recently, [ABF+25] presented a dramatically improved time–space tradeoff for SS-LE, affirmatively answering the first open question. They proposed a protocol parameterized by a positive integer $\rho$ ($1\leq\rho<n/2$) that solves SS-RK (and thus SS-LE) in $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ time w.h.p. using $O(\rho^{2}\log n)$ bits. Choosing $\rho=\omega(\log n)\cap o(\sqrt{n/\log n})$ yields a sublinear-time, sub-exponential-state protocol: for example, $\rho=\Theta(\sqrt{n}/\log n)$ gives $O(\sqrt{n}\log^{2}n)$ time w.h.p. with $2^{O(n/\log n)}=2^{o(n)}$ states, while $\rho=\Theta(\log^{2}n)$ gives $O(n/\log n)$ time w.h.p. with $2^{O(\log^{5}n)}$ states. Allowing super-exponentially many states—but still far fewer than in [BCC+21a]—yields a time-optimal solution: setting $\rho=\Theta(n)$ achieves $\Theta(\log n)$ time w.h.p. using $2^{O(n^{2}\log n)}$ states. Their collision-detection module also resolves the second question by [BCC+21a] independently of [AS25]. Although this module uses significantly more states ($2^{O(\rho^{2}\log n)}$ vs. $O(n\cdot\mathrm{poly}(\log n))$), it never terminates and thus does not require repeated invocation, enabling the first sublinear-time, sub-exponential-state SS-RK protocol.
@@
While the majority of work on population protocols studies the setting where every pair of agents can interact, quite a few studies deal with a more general setting: interactions may occur between restricted pairs of agents, introducing an interaction graph $G=(V,E)$, where $V$ represents the set of agents and $E$ the set of interactable pairs. SS-LE and its weaker variant have also been studied in rings [FJ06, CC19, YSM21, YSOM23], regular graphs [CC20], and general graphs [FJ06, SOK+18, SOKM20, SSN+21, KESI24].